Poisson probablity process with a known average of 6 cars per hour (per 60 minutes). a. Find the expected number of cars driving though in a 23 minutes period. (Round your final answer to1 decimal place.) b. Find the probability of at least 2 cars driving through in a given 23 minutes period. (Round your answer to 4 decimal places) Round you
media%2Faaf%2Faafd0330-eaf6-4ce3-a1e5-c3
Answer #1

As per HOMEWORKLIB POLICY, we are advised to do one question at a time so i am attemtping the second image.

For a continuous distibution, P(X = x) = 0

Hence,

a) P(X < 65) = 1 - P(X > 65) = 1 - P(X > 65) = 1 - 0.22 = 0.78

b) P(X < 21) = P(X < 65) - P(21 < X < 65) = 0.78 - 0.31 = 0.47

c) P(X = 65) = 0